# Decoding Reading Goals from Eye Movements

### v1

@@ 3.1 Eye Movements-only Models
Fixation-Based Representation
BEyeLSTM - No Text is our adaptation of the BEyeLSTM model (Reich et al., 2022) described below, which uses only eye movements input without the underlying text. This model is more expressive compared to the logistic regression model, in that rather than averaging eye movement measures for the entire trial, it explicitly represents each fixation in the eye movement trajectory. The final hidden state of an LSTM (see below), combined with global eye movement features, is then used for classification.
3.2 Eye Movements and Text Models
We examine a number of recent multimodal models that combine eye movements with textual information. The models encode textual information in two ways. The first is using contextual word embedding representations commonly used in NLP. The second is via linguistic word property features, including word length, word frequency and surprisal (Hale, 2001; Levy, 2008), which are motivated by their ubiquitous effects on reading times (Rayner et al., 2004; Kliegl et al., 2004; Rayner et al., 2011, among others).
The models implement three primary strategies for combining the two modalities at progressively later stages of processing: (i) in the model input, (ii) merging them within intermediate model representations, or (iii) with architectures that fuse the modalities using cross attention mechanisms after each modality has been processed separately. Furthermore, since eye movements in reading are both temporally and spatially aligned with the underlying text, the models can be categorized based on how they capture this alignment: (i) by aggregating eye movements for each word, thereby focusing on spatial alignment; or (ii) by aggregating eye movement information for each individual fixation, which explicitly encodes both spatial and temporal correspondences between eye movements and text.
@@ 3.2 Eye Movements and Text Models
PLM-AS (Yang & Hollenstein, 2023) This model represents the eye movements sequence by reordering contextual word embeddings according to the order of the fixations over the text. This reordered sequence is processed through a Recurrent Neural Network (RNN), whose final hidden layer is used for classification.
@@
BEyeLSTM (Reich et al., 2022) represents both the fixation sequence and textual features, combining LSTMs (Hochreiter & Schmidhuber, 1997) and global features through a linear layer.
@@
3.3 Baselines
When examining the utility of eye movements for a prediction task, it is important to benchmark models against simpler approaches that do not require eye movement information (Shubi et al., 2024). Here, we introduce two such baselines:
Reading Time (per word) Total reading time per word, computed by dividing the total reading time of the paragraph by the number of words in the paragraph. This behavioral baseline does not require eye-tracking and is motivated by the analyses of Malmaud et al. (2020), Hahn & Keller (2023) and Shubi & Berzak (2023), which indicate that on average, reading is faster in information seeking compared to ordinary reading.
4 Experimental Setup
4.1 Data
Addressing the proposed task is made possible by OneStop (Malmaud et al., 2020), the first dataset that contains broad coverage eye-tracking data in both ordinary reading and information seeking regimes. The textual materials of OneStop are taken from OneStopQA (Berzak et al., 2020), a multiple-choice reading comprehension dataset that comprises 30 Guardian articles from the OneStopEnglish corpus (Vajjala & Lučić, 2018). Each article is available in the original (Advanced) and simplified (Elementary) versions. Each paragraph has three multiple choice reading comprehension questions that can be answered based on any of the two paragraph difficulty level versions. Each question is paired with a manually annotated textual span, called the critical span, which contains the vital information for answering the question. The mean paragraph length is 109 words. An example of a OneStopQA paragraph along with one question and its critical span annotation is provided in Table 2 in Appendix A.
Eye movements data for OneStopQA were collected in-lab from 360 adult native English speakers using an EyeLink 1000+ eye tracker (SR Research), a state-of-the-art eyetracking device for reading research. Each participant read a batch of 10 articles (54 paragraphs) paragraph by paragraph and answered a comprehension question about each paragraph. Each paragraph was shown to a participant either in the Advanced or Elementary version, along with one of the three possible questions. The experiment has two between-subjects reading goal tasks: information seeking and ordinary reading. In the information seeking task, participants were presented with the question (without the answers) prior to reading the paragraph. In the ordinary reading task, they did not receive the question prior to reading the paragraph. In both tasks, after having read the paragraph, participants proceeded to answer the question on a new screen, without the ability to return to the paragraph.
In sum, each participant read 54 paragraphs, all of which are either in information seeking or in ordinary reading. Each paragraph was read by 120 participants: 60 in ordinary reading and 60 in information seeking (split equally between the Advanced and Elementary versions of the paragraph). Overall, the data consists of 19,438 trials, where a trial is a recording of eye movements from a single participant over a single paragraph. The data is balanced, with 9,718 trials in ordinary reading and 9,720 in information seeking. Figure 5 in Appendix A shows example trials for each of the two reading regimes. Additional data statistics are described in Appendix A.
4.2 Model Training and Evaluation Protocol
We adopt the training and evaluation procedure of Shubi et al. (2024) for prediction of reading comprehension from eye movements, with several modifications for our task. Similarly to Shubi et al. (2024), we use 10-fold cross validation, addressing three levels of model generalization:
  • New Item (paragraph): eye-tracking data is available during training for the participant but not for the paragraph.
  • New Participant: eye-tracking data is available during training for the paragraph, but not for the participant.
  • New Item & Participant: Neither the participant nor the paragraph are in the training data.
We further report aggregated results across all three regimes.
Figure 2: A schematic depiction of one of the 10 splits into train, validation, and the three test sets for one batch of 10 OneStopQA articles and 120 participants. Dashed lines denote information seeking trials. The full data split consists of the union of three such splits. Adapted from (Shubi et al., 2024).
The data is split separately for each batch of 10 articles and the 120 participants who read the batch. The three batch splits are then combined to form the full split of the dataset. Paragraphs are allocated to the train, validation and test portions of each batch split at the article level, such that all the paragraphs of each article appear in the same portion of the split. This ensures that items in the test set are unrelated in content to items in training and validation. Each split further guarantees an equal number of participants from each OneStopQA batch in each data portion.
Shubi et al. (2024) predicted reading comprehension separately in information seeking and ordinary reading, and thus defined separate data splits for the information seeking and ordinary reading trials. They further stratified items by the selected answers to the reading comprehension questions. As these steps are not relevant for our task, we use the union of the information seeking and ordinary reading trials in each part of each split, and stratify trials by reading goal rather than by reading comprehension answers.
Each data split contains approximately 64% of the trials in the training set, 17% in the validation set and 19% in the test sets (9% New Item, 9% New Participant and 1% New Item & Participant). Aggregated across the 10 splits, approximately 90% of the trials in the dataset appear in each of the New Participant and New Item evaluation regimes, and 10% in the New Item & Participant regime. Figure 2 presents this breakdown for one batch split.
We perform hyperparameter optimization and model selection separately for each split. We assume that at test time, the evaluation regime of the trial is unknown. Model selection is therefore based on the entire validation set of the split. All neural network-based models were trained using the PyTorch Lighting (Falcon & team, 2024) library on NVIDIA A100-40GB and L40S-48GB GPUs. Further details regarding the training procedure, including the full hyperparameter search space for all the models are provided in Appendix B.
4.3 Statistical Tests for Model Performance Comparisons
The samples in the OneStop dataset are not i.i.d; each item is read by multiple participants, and each participant reads multiple items. To account for these dependencies when performing statistical tests, we fit linear mixed-effects models with maximal random effects for items and participants (Barr et al., 2013). To this end, we use the MixedModels package in Julia (Bates et al., 2024).
5 Results
Test set accuracy results are presented in Table 1. In line with prior findings regarding faster reading in information seeking compared to ordinary reading (Malmaud et al., 2020; Hahn & Keller, 2023; Shubi & Berzak, 2023), the Reading Time baseline yields above chance accuracies ($p<0.01$ in all evaluation regimes), thus providing a strong benchmark for the eyetracking-based models. Among the 10 examined models, RoBERTa-Eye-F achieves the highest accuracy in the New Item, New Item & Participant and All evaluations, while RoBERTa-Eye-W slightly outperforms it in the New Participant regime. Differently from the New Item and the New Participant regimes, where multiple models outperform the Reading Time baseline, in the most challenging New Item & Participant regime, the improvements over this baseline are not statistically significant.
Table 1: Test accuracy results aggregated across 10 cross validation folds. ‘Emb’ stands for word embeddings, ‘Ling. Feat.’ for linguistic word properties such as word length, frequency and surprisal. Model performance is compared to the Reading Time baseline using a linear mixed effects model. In R notation: $is\_correct\sim model+(model\mid participant)+(model\mid paragraph)$. Significant gains over this baseline are marked with ’*’ $p<0.05$, ‘**’ $p<0.01$ and ‘***’ $p<0.001$.
Majority Class / Chance None None $50.0$ $50.0$ $50.0$ $50.0$
Reading Time None None $59.0$ $58.9$ $60.4$ $59.0$
[…]
MAG-Eye (Shubi et al., 2024) Words Emb + Ling. Feat. $52.1$ $52.3$ $51.5$ $52.1$
PLM-AS (Yang & Hollenstein, 2023) Fixations Order Emb $58.6$ $59.5$ $57.5$ $59.0$
Haller RNN (Haller et al., 2022) Fixations Emb $61.3$ $61.5$ $60.3$ $61.3$
[…]
Eyettention (Deng et al., 2023) Fixations Emb + Ling. Feat. $55.8$ $55.7$ $55.4$ $55.8$
[…]
In addition to RoBERTa-Eye-F, three other fixation based models, PostFusion-Eye and the two BEyeLSTM models, excel in the New Item regime. One possible explanation for this outcome is better generalization of fixation based models to new items compared to new participants. Note however, that due to the between-subjects design where all the training and test examples for a given participant have the same label, it could alternatively reflect, at least in part, an ability of fixation based models to learn participant specific reading behavior without explicit information on the participant (i.e. identify the participant), an ability that is not directly pertinent to the task at hand. The current experimental setup makes it challenging to adjudicate between these two possibilities. In either case, it is highly non-trivial that models are able to generalize from prior participant data to new items. In Appendix C, we further include ROC curves and AUC results (Bradley, 1997) in Figure 6, and validation set accuracies in Table 3. The outcomes of these evaluations are consistent with the test results reported in Table 1.
5.1 Logistic Ensemble
Ensemble models combine predictions from multiple individual models, and often yield better results than single models, especially when models have diverse predictive behaviors (Sagi & Rokach, 2018). To examine the appropriateness of this approach for our case, we compute the agreement between models’ predictions, quantified with Cohen’s Kappa (Cohen, 1960). Figure 3(a) depicts the pairwise agreement rates across models in the validation data, where we observe mostly moderate agreement rates. Figure 7 in Appendix D further suggests consistency across the three evaluation regimes. These outcomes indicate that while the examined models do not always differ substantially in accuracy, they likely capture different aspects of the eye movement data and the task.
Motivated by the results of this analysis, we introduce a simple Logistic Ensemble: a 10-feature logistic regression model that predicts the reading goal from the probability outputs of our 10 models. The model is trained for each fold on the validation set and tested on the corresponding three test sets. In Figure 3(b) we present the performance of the Logistic Ensemble, against the best performing model, RoBERTa-Eye-F, and the Reading Time baseline. We find that the Logistic Ensemble improves over the accuracy of RoBERTa-Eye-F in all the regimes, with an accuracy of 77.3 for New Items, 64.6 for New Participants, 64.3 for New Items & Participants, and 70.5 overall. The improvements over RoBERTa-Eye-F are statistically significant in all but the New Item & Participant regime, in which it is the only model that is statistically better than the Reading Time baseline. These performance improvements suggest that the information encoded by the different models is to some extent complementary.
Figure 3: (a) Pairwise Cohen’s Kappa agreement between model predictions on the validation set. (b) Test accuracy of the Logistic Ensemble model compared to the best single model RoBERTa-Eye-F and the Reading Time baseline for the ordinary reading vs. information-seeking goal decoding task, using 10-fold cross-validation. Statistical significance is tested with a linear mixed effects model. In R notation: $is\_correct\sim model+(model\mid participant)+(paragraph\mid parag)$. ’*’ $p<0.05$, ‘**’ $p<0.01$ and ‘***’ $p<0.001$.
6 What Makes the Task Easy or Hard?
Having established that the prediction task at hand can be performed with a considerable degree of success, we now leverage the best performing model to obtain insights about the task itself. We do so by examining which trial features, which were not given to the model explicitly, affect the ability of the model to classify trials correctly. This analysis takes advantage of the rich structure and auxiliary annotations of the OneStop dataset.
  1. 1-3.
    Reading time before, within, and after critical span: These features are motivated by the findings of Malmaud et al. (2020) and Shubi & Berzak (2023) regarding faster reading times in information seeking compared to ordinary reading, primarily before and after the critical span, as well as the reported classification results of the Reading Time baseline.
  2. 4.
    Paragraph position (1-54): Each participant reads 54 paragraphs, in a random article order. This feature captures the position of the paragraph in the experiment’s presentation sequence. It is included as reading strategies can change as the experiment progresses (e.g. Meiri & Berzak (2024) show that readers become faster as the experiment progresses).
  3. 5.
  1. 6.
  2. 7.
  3. 8.
    Critical span start location (relative position, normalized by paragraph length): Shubi & Berzak (2023) showed skimming-like reading patterns after processing task critical information in information seeking. We thus hypothesize that earlier appearance of task critical information could facilitate the ability to correctly identify information seeking reading.
  4. 9.
  5. 10.
    Question difficulty (percentage of participants who answered the question incorrectly): based on data from the Prolific QA experiment in Berzak et al. (2020). We include this feature as it can influence eye movements in information seeking, with harder questions potentially obscuring patterns of goal oriented reading.
In order to establish the relation of these features to task difficulty, we use a linear mixed effect model that uses these features to predict whether the trial prediction of the best performing model, RoBERTa-Eye-F, was correct. In R notation:
[…]
where the random effects account for correlations in predictions within participants, items and evaluation regimes [^2] Random effects structure is simplified not to include slopes due to model convergence issues.. The advantage of a multiple regression model over separate univariate analyses of each feature, is the ability to tease apart the contribution of each feature above and beyond all other features. We fit this model separately on the information seeking and ordinary reading trials. To make the contributions of the features to prediction accuracy comparable, we normalize each feature to be a z-score (zero mean and unit variance). We then examine feature contribution via statistical significance, magnitude and sign of the corresponding coefficient. A significant coefficient for a feature indicates that it correlates with task difficulty, the absolute value determines its importance rank relative to other features, and the sign indicates the direction of the association.
The resulting feature coefficients are presented in Figure 4. In line with the findings of Shubi & Berzak (2023), we observe that the most prominent features for correctly classifying both information seeking and ordinary reading trials are reading time before and after the critical span. Slower readers in these regions are easier to correctly classify as ordinary reading and harder to correctly classify as information seeking. We also find, as hypothesized, that shorter critical spans facilitate correct classification of information seeking trials, presumably by making information seeking more targeted. Finally, paragraph position is significant in information seeking, suggesting that readers develop more efficient goal oriented reading strategies as they progress through the experiment. Overall, this analysis provides a highly interpretable characterization of task difficulty, with insights on reading behavior in both reading regimes.
7 Related Work
The vast majority of the literature on the psychology of reading and its interfaces with computational modeling is concerned with ordinary reading. However, several studies did address goal-oriented (also referred to as task-based) reading. Most prior work focused on a small number of canonical tasks: skimming, speed reading and proofreading. Several studies found different eye movement patterns in these tasks as compared to ordinary reading (Just et al., 1982; Kaakinen & Hyönä, 2010; Schotter et al., 2014). Rayner & Raney (1996) examined differences between ordinary reading and searching through the text for a target word. Prior work also analyzed eye movements during human linguistic annotation, often used for generating training data for NLP tools, such as annotation of named entities (Tomanek et al., 2010; Tokunaga et al., 2017). Differences in reading patterns were also found when readers were asked to take different perspectives on a given text (Kaakinen et al., 2002) or given different sets of learning goals (Rothkopf & Billington, 1979).
Our work is closest to Malmaud et al. (2020), Hahn & Keller (2023) and Shubi & Berzak (2023) who analyzed eye movement differences between ordinary reading and information seeking, where the information seeking goal is formulated using a reading comprehension question. All three studies found substantial differences in fixation and saccade patterns in information seeking as compared to ordinary reading, in particular before, within and after the text portions that are critical for the information seeking task. Here, we build on these findings, and examine whether these differences can be leveraged for automatic discrimination between these two reading regimes.
While the above studies focus primarily on descriptive data analysis, Hollenstein et al. (2023) took a predictive approach and attempted to automatically classify the reading task from eye movement features. In this study, participants read single sentences from the ZuCo corpus (Hollenstein et al., 2020), and engaged either in ordinary reading or in an annotation task where they had to determine the presence of one of seven semantic relations in the sentence. While this benchmark is conceptually related to the current work, it is restricted in scope and ecological validity. First, the benchmark contains only 18 participants, which makes generalization and reliable evaluations highly challenging. More fundamentally, the framework is limited by the nature of the tasks, which focus on highly specialized linguistic annotations that are not performed by readers in everyday life. If viewed through the lens of question answering, the setup comprises only seven possible questions that concern the presence of a semantic relation in the sentence. In the current study we take a different and more general stance on task based reading, with unrestricted questions that are more representative of the tasks commonly pursued by readers.
8 Summary and Discussion
Is it possible to decode reader goals from eye movements? We address this question by examining the possibility to automatically discriminate between ordinary reading and information seeking at the challenging granularity level of a single paragraph. We find that it is indeed possible to perform this task with considerable success, and show that eye movements provide valuable information for this task, above and beyond reading time. We further find that different models capture different aspects of the signal and the task, and can be successfully combined with ensembling, which yields the strongest results for our task. Our error analysis leverages the models to further reveal new insights on the factors that determine task difficulty.
While the current work takes a first step in addressing the proposed task, ample room for performance improvements remains for future work, especially in the new participant regimes. New strategies for modeling eye movements with text are likely needed to fully exploit the potential of eye movements for this task. Furthermore, the addressed task is fundamentally limited in that it does not distinguish between different information seeking tasks. A natural direction for future work could address decoding of the specific question that was presented to the participant in the information seeking regime. Finally, new datasets for the information seeking task, other types of tasks, additional populations (e.g. second language readers, younger and older participants), and datasets in languages other than English are all needed in order to study goal decoding more broadly.
9 Ethics Statement
This work uses eye movement data collected from human participants. The data was collected by Malmaud et al. (2020) under an institutional IRB protocol. All the participants provided written consent prior to participating in the eyetracking study. The data is anonymized.
10 Reproducibility Statement
We describe the model training and selection procedure, the evaluation protocol, and the hardware and software specifications in Section 4.2. The hyperparameter search space for each model is described in Appendix B. The dataset used in the experiments is described in Sections 4.1 and A. The source code, which implements the models, experimental procedures, and analysis, is available at: https://github.com/lacclab/Goal-Decoding-from-Eye-Movements. Data will be made publicly available.
Appendix
Appendix A OneStop Eye Movements Dataset - Additional Details
The textual data of OneStop consists of $162$ paragraphs, $486$ questions, and $972$ unique paragraph–level–question triplets. The mean length of Elementary paragraphs is 97 words (37 before the critical span, 30 inside it, and 30 after it), and of Advanced paragraphs 120 words (48 before the critical span, 34 inside it, and 38 after it). Each question has 20 responses, 10 for the Advanced version and 10 for the Elementary version. The mean experiment duration is approximately one hour. The raw millisecond gaze location is pre-processed into fixations and saccades using the SR Data Viewer software (v4.3.210).
@@ D Negotiating extradition agreements for drug traffickers
Appendix B Model Training and Hyperparameters
@@
  • Following (Reich et al., 2022), for BEyeLSTM and BEyeLSTM-No Text, the search space consists of learning rates $\{0.001,0.003,0.01\}$, embedding dimensions $\{4,8\}$ and hidden dimensions $\{64,128\}$.
@@
  • Following Yang & Hollenstein (2023), we train PLM-AS and Haller RNN with dropout rate search space of 0.1, and for PLM-AS, we use LSTM layer counts of ${1,2}$. Additionally, as in (Haller et al., 2022), we search over LSTM hidden layer sizes of ${10,40,70}$. For PLM-AS, the LSTM hidden layer size is fixed at 1024 to match the LM’s dimensionality in Yang & Hollenstein (2023).
@@
We train the deep-learning based models for a maximum of 40 epochs, with early stopping after 8 epochs if no improvement in the validation error is observed. Following Liu et al. (2019); Mosbach et al. (2021); Shubi et al. (2024) we use the AdamW optimizer Loshchilov & Hutter (2018) with a batch size of $16$. MAG-Eye, RoBERTa-Eye and PostFusion-Eye use a linear warm-up ratio of $0.06$, and a weight decay of 0.1. We standardize each eye movements feature using statistics computed on the training set, to zero mean unit variance. Models are implemented based on code from (Shubi et al., 2024).
Appendix C Additional Results
Below we present test set ROC curves and AUC. We also provide accuracy results for the validation set.
Figure 6: ROC Curves by model and evaluation regime. Each curve represents a different model, with the corresponding Area Under the Curve (AUC) values provided in the legend.
Table 3: Validation accuracy results aggregated across 10 cross-validation folds. ’Emb’ stands for word embeddings, ’Ling. Feat.’ for linguistic word properties. Model performance is compared to the Reading Time baseline using a linear mixed-effects model. In R notation: $is\_correct\sim model+(model\mid participant)+(model\mid paragraph)$. Significant gains over this baseline are marked with ’*’ $p<0.05$, ’**’ $p<0.01$, and ’***’ $p<0.001$.

### v3

@@ A.1 Model Descriptions
PLM-AS (Yang and Hollenstein, 2023) This model represents the eye movements sequence by reordering contextual word embeddings according to the order of the fixations over the text. This reordered sequence is processed through a Recurrent Neural Network (RNN), whose final hidden layer is used for classification.
@@
BEyeLSTM (Reich et al., 2022a) represents both the fixation sequence and textual features, combining LSTMs (Hochreiter and Schmidhuber, 1997) and global features through a linear layer.
BEyeLSTM - No Text is a model that processes raw fixation data using an LSTM. The final hidden state of the LSTM is combined with global eye movement features to perform classification. The model is inspired by BEyeLSTM (Reich et al., 2022a), using the same eye movements feature set, without the text representations.
@@
Image Representations
We represent scanpaths as two-dimensional images, as illustrated in Figure 4. In this visualization, fixations are depicted as circles positioned at their original x-y coordinates from the screen display, with the entire representation cropped to maintain consistent dimensions. The diameter of each circle corresponds to the duration of the fixation. To indicate the sequential progression of the eye movements, we employ a gradient shading scheme. Additionally, we differentiate between saccade types by color-coding them according to the five categories established by Schotter and Dillon (2025) - forward saccade, skip, refixation, return sweep, regression, and another for any saccade typethat does not fall into this categorization. Note that for these features knowledge about the existance of text is needed, but not the textual content itself. We use the convnextv2_base.fcmae_ft_in22k_in1k and vit_base_patch14_dinov2 versions of the ConvNextv2 and ViT models respectively.
Figure 4: An example of a scanpath as an image as used for the image classification models.
A.2 Model Diagrams
(a) BEyeLSTM - No Text
(b) BEyeLSTM
(c) Eyettention
(d) Haller RNN
Figure 5: Visualization of the different model architectures (Part 1). $P$ represents the paragraph, $E^{P}_{S}$ the eye movements of participant $S$ on $P$. $LM$ stands for a language model, and $FC$ for fully connected layers. $FF_{f_{i}}$ stands for the fixation features and $w_{f_{i}}$ for the word corresponding to the $i$-th fixation respectively.
(a) PostFusion-QEye
(b) PLM-AS
(c) MAG-QEye
(d) RoBERTa-QEye
Figure 6: Visualization of the different model architectures (Part 2). $P$ represents the paragraph, $E^{P}_{S}$ the eye movements of participant $S$ on $P$. $LM$ stands for a language model, and $FC$ for fully connected layers. $FF_{f_{i}}$ stands for the fixation features and $w_{f_{i}}$ for the word corresponding to the $i$-th fixation respectively.
Appendix B OneStop Eye Movements Dataset - Additional Details
The textual data of OneStop consists of $162$ paragraphs, $486$ questions, and $972$ unique paragraph–level–question triplets. The mean paragraph length is 109 words (min: 50; max: 165; std: 28). The mean length of Elementary paragraphs is 97 words (37 before the critical span, 30 inside it, and 30 after it), and of Advanced paragraphs 120 words (48 before the critical span, 34 inside it, and 38 after it). Each question has 20 responses, 10 for the Advanced version and 10 for the Elementary version. The mean experiment duration is approximately one hour. The raw millisecond gaze location is pre-processed into fixations and saccades using the SR Data Viewer software (v4.3.210).
@@ D Negotiating extradition agreements for drug traffickers
Appendix C Model Training and Hyperparameters
All neural network-based models were trained using the PyTorch Lighting (Falcon and team, 2024) library on NVIDIA A100-40GB and L40S-48GB GPUs.
@@ Appendix C Model Training and Hyperparameters
  • Following (Reich et al., 2022a), for BEyeLSTM and BEyeLSTM - No Text, the search space consists of learning rates $\{0.001,0.003,0.01\}$, embedding dimensions $\{4,8\}$ and hidden dimensions $\{64,128\}$.
@@
  • Following Yang and Hollenstein (2023), we train PLM-AS and Haller RNN with dropout rate search space of 0.1, and for PLM-AS, we use LSTM layer counts of ${1,2}$. Additionally, as in (Haller et al., 2022), we search over LSTM hidden layer sizes of ${10,40,70}$. For PLM-AS, the LSTM hidden layer size is fixed at 1024 to match the LM’s dimensionality in Yang and Hollenstein (2023).
@@
We train the deep-learning based models for a maximum of 40 epochs, with early stopping after 8 epochs if no improvement in the validation error is observed. Following Liu et al. (2019); Mosbach et al. (2021); Shubi et al. (2024) we use the AdamW optimizer Loshchilov and Hutter (2018) with a batch size of $16$. MAG-Eye, RoBERTa-Eye and PostFusion-Eye use a linear warm-up ratio of $0.06$, and a weight decay of 0.1. We standardize each eye movements feature using statistics computed on the training set, to zero mean unit variance.
The code base for this project was developed with the assistance of GitHub Copilot, an AI-powered coding assistant. All generated code was carefully reviewed.
Appendix D Additional Results
Below we present the test set ROC curves across the ten cross-validation splits and their corresponding AUROC scores (mean and standard deviation). We also provide accuracy results for the validation set.
Figure 8: ROC Curves by model and evaluation regime. Each curve represents a different model across the ten cross-validation splits, with the corresponding AUROC scores (mean and standard deviation) provided in the legend.